mhchen:

Prove that there exists a real number 'x' satisfying x^2 = 2

mhchen:

\[T = \left\{ t \in R : t^{2} < 2\right\}\] set x = least upper bound of T. We need to prove x^2 = 2

mhchen:

Let supT = "least upper bound of T" To prove this, we shall consider this: If a and b are real numbers, then either a < b or a = b or a > b. To prove that x^2 = 2 is supT We need to prove that x^2 < 2 is false, and x^2 > 2 is false.

mhchen:

Case 1: Prove that x^2 < 2 is false by proof by contradiction. Assume x^2 < 2 Then \[(x + \frac{1}{n})^{2} < 2\] Is not an element of set T

mhchen:

\[(x + \frac{1}{n})^{2} < 2\] then \[=x^2 + \frac{2x}{n}+\frac{1}{n} < x^{2} + \frac{2x}{n}+\frac{1}{n} = x^{2}+\frac{2x+1}{n}\] By the archimedean property, for any real number x, there exists an n such that: \[\frac{1}{n} < x \]

mhchen:

But the archimedean property is useless here because I'm trying to find an 'n' that makes the previous equation greater than x^2 and less than 2. Instead, choose n such that n is a natural number, and \[\frac{1}{n} < \frac{2 - x^{2}}{2x+1}\]

mhchen:

Then \[\frac{2x + 1}{n} < 2 - x^{2}\] then \[(x + \frac{1}{n})^{2} < x^{2} + (2 - x^{2}) = 2\] then (x+1)/n is an element of T. Contradicting the fact that x is an upper bound of T. Thus x^2 < 2 is false.

mhchen:

Case 2: Prove the same thing about x^2 > 2.